OpenStudy (anonymous):

Compact fluorescent bulbs are much more efficient at producing light than are ordinary incandescent bulbs. They initially cost much more, but last far longer and use much less electricity. According to one study of these bulbs, a compact bulb that produces as much light as a 100 {\rm W} incandescent bulb uses only 23.0 {\rm W} of power. The compact bulb lasts 1.00×104 hours, on the average, and costs \$ 11.0 , whereas the incandescent bulb costs only 76.0 ¢, but lasts just 750 hours. The study assumed that electricity cost 8.00 ¢ per {\rm kWh} and that the bulbs were on for 4.0 {\rm h} per day

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

Part B What is the total cost (including the price of the bulbs) to run compact fluorescent bulbs for 3.0 years? ________ dollar(s) Part C How much do you save over 3.0 years if you use a compact fluorescent bulb instead of an incandescent bulb? ________ dollar(s)

OpenStudy (anonymous):

To find the cost of operation. First, we need to find the number of hours of operation will occur per year. \[t_{op} = 4 \left [\rm hr \over day \right] \cdot \left [ \rm 365 ~day \over year \right ] \cdot \left [ \rm 3~ years \right] \]Now we need to find the number of bulbs that will be needed. \[N_{bulb} = {t_{op} \over t_{life}}\]Now we can find the cost. \[Cost = C_{kWh} \cdot t_{op} \cdot W + C_{bulb} \cdot N_{bulb}\]where \(C_{kWh}\) is the price per kWh and \(C_{bulb}\) is the cost per bulb. Find both costs and take the difference for Part C.

OpenStudy (anonymous):

Negative. This is YOUR homework. YOU have to run the numbers.